# Invariants of persistence modules defined by order-embeddings

### v2

@@ 4.1. Exact structures and relative projectives
[…]
@@
    $\mathcal{E}$ is closed under compositions of admissible monomorphisms, that is, if $f:X\to Y$ and $g:Y\to Z$ are admissible monomorphisms, then $g\circ f:X\to Z$ is also an admissibe monomorphism. Likewise, $\mathcal{E}$ is closed under compositions of admissible epimorphisms.
@@
We verify the three points from definition 4.1 for all $f\in\mathrm{E}$, using that the restriction functor $f^{*}:\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ is additive and exact.
@@
$[L]+[N]-[M]\textrm{ for any short exact sequence }(0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0)\textrm{ in }\mathcal{E}.$
The following is then immediate.
@@
Let $\mathcal{X}$ be a representation-finite poset and $\mathcal{P}$ be a finite poset. Recall that $\mathrm{I}=\operatorname{\mathrm{Ind}}\mathcal{X}$ and fix $\mathrm{E}\subseteq\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Then denote by $\mathcal{E}^{\mathrm{E}}_{\mathcal{P}}:=\Psi_{\mathrm{E}}(\mathcal{E}^{\rm sp}_{\mathcal{X}})$, where $\mathcal{E}_{\mathcal{X}}^{\rm sp}$ is the split exact structure on $\operatorname{\mathrm{mod}}\mathcal{X}$. Then we have a factorization
[…]
@@ 4.2. Relative projectives
    We say an object $P\in\operatorname{\mathrm{mod}}\mathcal{P}$ is $\mathcal{E}$-projective, in symbols $P\in\operatorname{\mathrm{Proj}}(\mathcal{E})$, if $g_{*}=\operatorname{\mathrm{Hom}}_{\mathcal{P}}(P,g)$ is an epimorphism for all $Y,Z\in\operatorname{\mathrm{mod}}\mathcal{P}$ and for all admissible epimorphisms $g:Y\to Z$. Equivalently, $P\in\operatorname{\mathrm{Proj}}(\mathcal{E})$ if and only if $\operatorname{\mathrm{Hom}}_{\mathcal{A}}(P,-)$ sends admissible exact sequences to exact sequences.
@@
    We say an object $I\in\operatorname{\mathrm{mod}}\mathcal{P}$ is $\mathcal{E}$-injective, in symbols $I\in\operatorname{\mathrm{inj}}(\mathcal{E})$, if $f_{*}=\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f,I)$ is an epimorphism for all $X,Y\in\operatorname{\mathrm{mod}}\mathcal{P}$ and for all admissible monomorphisms $f:X\to Y$. Equivalently, $I\in\operatorname{\mathrm{inj}}(\mathcal{E})$ if and only if $\operatorname{\mathrm{Hom}}_{\mathcal{P}}(-,I)$ sends admissible exact sequences to exact sequences.
@@
[…]
@@
Let $f:\mathcal{X}\to\mathcal{P}$ be an order embedding. Then the restriction functor $f^{*}:\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ is exact and has a left adjoint $f_{!}:\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$ and a right adjoint $f_{*}:\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$. Moreover we have $f^{*}f_{!}={\rm Id}_{\operatorname{\mathrm{mod}}\mathcal{X}}=f^{*}f_{*}$.
The functors $f^{*}$ and $f_{!}$ are sometimes referred to as the induction and coinduction functors, respectively. The induction functor preserves projective presentations and the coinduction functor preserves injective copresentations.
@@
$\mathcal{H}^{\mathrm{E}}:=\{f_{!}U\mid(f,U)\in\mathrm{E}\times\mathrm{I}\}.$
@@
First note that since $\mathcal{P}=\bigcup_{f\in\mathrm{E}}\operatorname{\mathrm{Im}}f$, for any $p\in\mathcal{P}$ there exists $x\in\mathcal{X}$ such that $f(x)=p$. Since the functor $f_{!}$ sends projectives to projectives, we have that $f_{!}(P_{x})=P_{p}$, that is the projective $P_{p}$ is in $\mathcal{H}^{\mathrm{E}}$. Moreover by the adjunction formula we have for any $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$
@@
which is local since $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$. Note that since $\mathcal{X}$ is representation-finite, the module $U$ and hence the module $f_{!}U$ is even a brick, that is a module whose endomorphism is one dimensional.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$(f,U)\sim_{!}(g,V)\quad\Leftrightarrow\quad f_{!}U\simeq g_{!}V.$
@@
    $\{f_{!}U\mid U\in I\}=\left\{P_{f1},P_{f2},P_{f3},H_{f1,f2},H_{f1,f3},M:=\operatorname{\mathrm{Coker}}(P_{f2}\oplus P_{f3}\to P_{f1})\right\}.$
@@
    $f(1)$$f(2)$$f(3)$$k$
@@
    $\dim M_{x}=\left\{\begin{array}[]{ll}k&\textrm{if }x\geq f(1)\\ k&\textrm{if }x\geq f(2)\textrm{ and }x\ngeq f(3)\\ k&\textrm{if }x\geq f(3)\textrm{ and }x\ngeq f(2)\\ k^{2}&\textrm{if }x\geq f(2),\ x\geq f(3)\textrm{ and }x\ngeq f(1)\\ 0&\textrm{else},\end{array}\right.$
@@
    $f(3)$$f(2)$$f(1)$$k$$k$$k$$k^{2}$
@@ 4.3. Bases for invariants and relative projectives
Let $\Psi:\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{N}$ be a finite invariant. A basis of $\Psi$ is a set $\mathcal{B}$ of indecomposable $\mathcal{P}$-modules such that $\Psi$ induces an isomorphism of abelian group
@@
Let $\mathcal{X}$ be a representation-finite poset and $\mathcal{P}$ be a finite poset. Recall that $\mathrm{I}=\operatorname{\mathrm{Ind}}\mathcal{X}$ and fix $\mathrm{E}$ a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Assume that $\bigcup_{f\in\mathrm{E}}\operatorname{\mathrm{Im}}f=\mathcal{P}$ and that the Gabriel quiver of the endomorphism algebra $\operatorname{\mathrm{End}}_{\mathcal{P}}(\bigoplus_{(f,U)\in\mathrm{E}\times\mathrm{I}/\sim_{!}}f_{!}U)$ has no oriented cycle. Then the set $\mathcal{H}_{\mathrm{E}}:=\{f_{!}U\mid(f,U)\in\mathrm{E}\times\mathrm{I}/{\sim_{!}}\}$ is a $\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$-basis (and hence a $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$-basis).
@@
[…]
@@ 5.1. Interval modules and the intermediate extension functor
We highlight that Example 4.12 gives two distinct bases for the invariants $\operatorname{\mathbf{dim}}$ and $\operatorname{\mathbf{rk}}$. In each case, one basis comes directly from the interpretation of the relevant invariant as a dim-hom invariant, while the other is somewhat less clear from a homological perspective. In this case, we use poset embeddings to give a unified homological interpretation of the fact that the simple modules form a basis of $\operatorname{\mathbf{dim}}$ and the fact that the rectangle modules form a basis of $\operatorname{\mathbf{rk}}$. In doing so, we also uncover new bases of several other invariants.
@@
${(3,1)}$${(3,2)}$${(3,3)}$${c}$${(2,1)}$${(2,2)}$${(2,3)}$${a}$${d}$${(1,1)}$${(1,2)}$${(1,3)}$${b}$${\mathcal{P}}$${\mathcal{X}}$
Let $f:\mathcal{X}\rightarrow\mathcal{P}$ be the poset embedding given by $fa=(2,1)$, $fb=(1,2)$, $fc=(3,2)$, and $fd=(2,3)$. Consider the sincere interval module $\mathbb{I}_{[\mathcal{X}]}\in\operatorname{\mathrm{mod}}\mathcal{X}$. Since the induction functor $f_{!}$ preserves projective presentations and the coinduction functor $f_{*}$ preserves injective copresentations, we then have
@@
where $\iota_{i}$ denotes the $i$-th inclusion map, $\mathrm{pr}_{i}$ denotes the $i$-th projection map, and $\Delta$ and $\nabla$ denote the diagonal and co-diagonal maps, respectively. We note that both $f_{!}\mathbb{I}_{[\mathcal{X}]}$ and $f_{*}\mathbb{I}_{[\mathcal{X}]}$ are indecomposable but are not interval modules. Moreover, we have that $\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}\mathbb{I}_{[\mathcal{X}]},f_{*}\mathbb{I}_{[\mathcal{X}]})\cong K$ and that the image of any nonzero map $f_{!}\mathbb{I}_{[\mathcal{X}]}\rightarrow f_{*}\mathbb{I}_{[\mathcal{X}]}$ is precisely $\mathbb{I}_{S}$ for $S=\{(i,j)\in\mathcal{P}\mid 3\leq i+j\leq 5\}$.
@@
Let $S\subseteq\mathcal{P}$. The convex hull of $\mathcal{P}$ is the set
@@
Let $f:\mathcal{X}\rightarrow\mathcal{P}$ be a poset embedding and let $U\in\operatorname{\mathrm{mod}}\mathcal{X}$. By adjunction and the fact that $f^{*}f_{*}=\mathrm{Id}$, there is a natural isomorphism $\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}U,f_{*}U)\cong\operatorname{\mathrm{End}}_{\mathcal{X}}(U)$. Let $\theta_{f}U:f_{!}U\rightarrow f_{*}U$ denote the image of the identity map $1_{U}\in\operatorname{\mathrm{End}}_{\mathcal{X}}(U)$ under this isomorphism. We denote $\Theta_{f}U=\mathrm{Im}(\theta_{f}U)$.
@@
[AET25, Proposition 4.16] Let $f:\mathcal{X}\rightarrow\mathcal{P}$ be a poset embedding and $\mathbb{I}_{S}\in\operatorname{\mathrm{mod}}\mathcal{X}$ an interval module. Then $\Theta_{f}(\mathbb{I}_{S})\cong\mathbb{I}_{\overline{fS}}$.
The next example uses Proposition 5.5 to explain the relationship between the hook modules and rectangle modules in studying the rank invariant.
@@
    Let $\mathcal{X}_{2}=\{1\rightarrow 2\}$. We recall from Example 3.7 that the three indecomposable $\mathcal{X}_{2}$-modules are $\mathbb{I}_{\{1,2\}}=P_{1}=I_{2}$, $\mathbb{I}_{\{1\}}=H_{1,2}=I_{a}$, and $\mathbb{I}_{\{2\}}=P_{2}=C_{1,2}$. For $f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$, we then compute
@@
    In particular, we have that $\{\Theta_{f}U\mid f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}},U\in\operatorname{\mathrm{Ind}}\mathcal{X}\}$ is the set of rectangle modules. Note that there is some redundancy in this collection: for $x\in\mathcal{P}$ we obtain the (simple) rectangle module $\mathbb{I}_{\{x\}}$ once for every embedding $f:\mathcal{X}_{2}\rightarrow\mathcal{P}$ with either $x=f(1)$ or $x=f(2)$.
@@
    but the invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ does not factor through $\mathbb{Z}^{\mathrm{E}\times\mathrm{I}/\sim_{\Theta}}$. For example, if $\mathcal{P}=\{a\to b\to c\}$, then we have $\Theta_{f}\mathbb{I}_{\{1\}}=\mathbb{I}_{\{a\}}=\Theta_{g}\mathbb{I}_{\{1\}}$ for $\operatorname{\mathrm{Im}}f=\{a,b\}$ and $\operatorname{\mathrm{Im}}g=\{a,c\}$. However, we have for any $\mathcal{P}$-module $M$ such that $\operatorname{\mathrm{rk}}M_{a\to b}\neq\operatorname{\mathrm{rk}}M_{a\to c}$ that

### v3

@@ 4.1. Exact structures and relative projectives
Let $\mathcal{E}$ be a class of short exact sequences in $\operatorname{\mathrm{mod}}\mathcal{P}$. We assume that $\mathcal{E}$ is closed under isomorphisms. If a short exact sequence $\eta=(0\to A\xrightarrow{f}E\xrightarrow{g}B\to 0)$ belongs to $\mathcal{E}$, we say that $\eta$ is an ($\mathcal{E}$-)admissible short exact sequence, and also that $f$ is an ($\mathcal{E}$-)admissible monomorphism and $g$ is an ($\mathcal{E}$-)admissible epimorphism. The closure under isomorphisms then implies that the classes of admissible short exact sequences, admissible monomorphisms, and admissible epimorphisms uniquely determine one another. Quillen’s definition can then be rephrased as follows:
@@
    $\mathcal{E}$ is closed under compositions of admissible monomorphisms, that is, if $f\mathrel{\mathop{\ordinarycolon}}X\to Y$ and $g\mathrel{\mathop{\ordinarycolon}}Y\to Z$ are admissible monomorphisms, then $g\circ f\mathrel{\mathop{\ordinarycolon}}X\to Z$ is also an admissibe monomorphism. Likewise, $\mathcal{E}$ is closed under compositions of admissible epimorphisms.
@@
We verify the three points from definition 4.1 for all $f\in\mathrm{E}$, using that the restriction functor $f^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ is additive and exact.
@@
[…]
@@
Let $\mathcal{X}$ be a representation-finite poset and $\mathcal{P}$ be a finite poset. Recall that $\mathrm{I}=\operatorname{\mathrm{Ind}}\mathcal{X}$ and fix $\mathrm{E}\subseteq\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Then denote by $\mathcal{E}^{\mathrm{E}}_{\mathcal{P}}\mathrel{\mathop{\ordinarycolon}}=\Psi_{\mathrm{E}}(\mathcal{E}^{\rm sp}_{\mathcal{X}})$, where $\mathcal{E}_{\mathcal{X}}^{\rm sp}$ is the split exact structure on $\operatorname{\mathrm{mod}}\mathcal{X}$. Then we have a factorization of the commutative square given in Theorem 3.6
${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}}$${\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${{\rm K}_{0}(\mathcal{E}_{\mathcal{P}}^{\mathrm{E}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}}$${\mathbb{Z}^{\mathrm{E}\times\mathrm{I}}}$
Proof.
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$ be in $\mathrm{E}$ and $(0\rightarrow L\rightarrow M\rightarrow N\rightarrow 0)$ be an $\mathcal{E}$-exact sequence. Then by definition we have an isomorphism $f^{*}M\simeq f^{*}N\oplus f^{*}L$. Therefore for each $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$ we have
$\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}(M)_{(f,U)}\mathrel{\mathop{\ordinarycolon}}=\mathbf{mult}_{\mathcal{X}}(f^{*}M)_{U}=\mathbf{mult}_{\mathcal{X}}(f^{*}N)_{U}+\mathbf{mult}_{\mathcal{X}}(f^{*}L)_{U},$
which proves the desired factorization. ∎
@@ 4.2. Relative projectives
    We say an object $P\in\operatorname{\mathrm{mod}}\mathcal{P}$ is $\mathcal{E}$-projective, in symbols $P\in\operatorname{\mathrm{Proj}}(\mathcal{E})$, if $g_{*}=\operatorname{\mathrm{Hom}}_{\mathcal{P}}(P,g)$ is an epimorphism for all $Y,Z\in\operatorname{\mathrm{mod}}\mathcal{P}$ and for all admissible epimorphisms $g\mathrel{\mathop{\ordinarycolon}}Y\to Z$. Equivalently, $P\in\operatorname{\mathrm{Proj}}(\mathcal{E})$ if and only if $\operatorname{\mathrm{Hom}}_{\mathcal{A}}(P,-)$ sends admissible exact sequences to exact sequences.
@@
    We say an object $I\in\operatorname{\mathrm{mod}}\mathcal{P}$ is $\mathcal{E}$-injective, in symbols $I\in\operatorname{\mathrm{inj}}(\mathcal{E})$, if $f_{*}=\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f,I)$ is an epimorphism for all $X,Y\in\operatorname{\mathrm{mod}}\mathcal{P}$ and for all admissible monomorphisms $f\mathrel{\mathop{\ordinarycolon}}X\to Y$. Equivalently, $I\in\operatorname{\mathrm{inj}}(\mathcal{E})$ if and only if $\operatorname{\mathrm{Hom}}_{\mathcal{P}}(-,I)$ sends admissible exact sequences to exact sequences.
@@
$\mathcal{E}_{\mathcal{U}}\mathrel{\mathop{\ordinarycolon}}=\left\{(0\to M\to N\overset{g}{\to}L\to 0)\mid\forall U\in\mathcal{U},\ \operatorname{\mathrm{Hom}}_{\mathcal{P}}(U,g)\textrm{ is surjective.}\right\}$
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$ be an order embedding. Then the restriction functor $f^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ is exact and has a left adjoint $f_{!}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$ and a right adjoint $f_{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$. Moreover we have $f^{*}f_{!}={\rm Id}_{\operatorname{\mathrm{mod}}\mathcal{X}}=f^{*}f_{*}$.
The functors $f_{*}$ and $f_{!}$ are sometimes referred to as the induction and coinduction functors, or right and left Kan extensions, respectively. The induction functor preserves projective presentations and the coinduction functor preserves injective copresentations.
@@
$\mathcal{H}^{\mathrm{E}}\mathrel{\mathop{\ordinarycolon}}=\{f_{!}U\mid(f,U)\in\mathrm{E}\times\mathrm{I}\}.$
@@
First note that since $\mathcal{P}=\bigcup_{f\in\mathrm{E}}\operatorname{\mathrm{Im}}f$, for any $p\in\mathcal{P}$ there exists $x\in\mathcal{X}$ and $f\in\mathrm{E}$ such that $f(x)=p$. Since the functor $f_{!}$ sends projectives to projectives, we have that $f_{!}(P_{x})=P_{p}$, that is the projective $P_{p}$ is in $\mathcal{H}^{\mathrm{E}}$. Moreover by the adjunction formula we have for any $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$
@@
which is local since $U\in\operatorname{\mathrm{Ind}}\mathcal{X}$.
@@
${\epsilon=(0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0)}$
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${f^{*}L\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${f^{*}M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${f^{*}N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{X}}(U,f^{*}L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{X}}(U,f^{*}M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{X}}(U,f^{*}N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
@@
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}U,L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}U,M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}U,N)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
@@
$(f,U)\sim_{!}(g,V)\quad\Leftrightarrow\quad f_{!}U\simeq g_{!}V$
and the map $\mathbb{Z}^{\mathcal{H}^{\mathrm{E}}}\to\mathbb{Z}^{\mathrm{E}\times\mathrm{I}/\sim_{!}}$ is induced by the map $\mathrm{E}\times\mathrm{I}\to\mathcal{H}^{\mathrm{E}}$ defined by $(f,U)\mapsto f_{!}U$.
@@
    $\{f_{!}U\mid U\in I\}=\left\{P_{f1},P_{f2},P_{f3},H_{f1,f2},H_{f1,f3},M\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathrm{Coker}}(P_{f2}\oplus P_{f3}\to P_{f1})\right\}.$
@@
    In the case where $\mathcal{P}$ is a grid poset, the $\mathcal{P}$-module $M$ looks as follows.
    f(1)f(1)f(2)f(2)f(3)f(3)$k$
@@
[…]
@@
    In the case where $\mathcal{P}$ is a grid poset, the $\mathcal{P}$-module $M$ looks as follows.
[…]
@@ 4.3. Bases for invariants and relative projectives
Let $\Psi\mathrel{\mathop{\ordinarycolon}}\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\to\mathbb{Z}^{N}$ be a finite invariant. A basis of $\Psi$ is a set $\mathcal{B}$ of indecomposable $\mathcal{P}$-modules such that $\Psi$ induces an isomorphism of abelian group
@@
Let $\mathcal{X}$ be a representation-finite poset and $\mathcal{P}$ be a finite poset. Recall that $\mathrm{I}=\operatorname{\mathrm{Ind}}\mathcal{X}$ and fix $\mathrm{E}$ a subset of $\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$. Assume that $\bigcup_{f\in\mathrm{E}}\operatorname{\mathrm{Im}}f=\mathcal{P}$ and that the Gabriel quiver of the endomorphism algebra $\operatorname{\mathrm{End}}_{\mathcal{P}}(\bigoplus_{(f,U)\in\mathrm{E}\times\mathrm{I}/\sim_{!}}f_{!}U)$ has no oriented cycle. Then the set $\mathcal{H}^{\mathrm{E}}\mathrel{\mathop{\ordinarycolon}}=\{f_{!}U\mid(f,U)\in\mathrm{E}\times\mathrm{I}/{\sim_{!}}\}$ is a $\mathbf{dimh}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$-basis (and hence a $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{\mathrm{E}}$-basis).
@@
${P_{a_{i}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P_{b_{i}}\oplus P_{c_{i}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P_{a_{i+1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P_{b_{i+1}}\oplus P_{c_{i+1}}}$
@@ 5. Interval bases via families of embeddings
In this section, we first reinterpret classical bases for invariants, such as the simple modules for the dimension vector and the rectangle modules for the rank invariant, in terms of the intermediate extension functor. This provides a conceptual explanation for their appearance.
We then develop a more general framework based on families of embeddings of representation-finite posets, which allows us to systematically construct bases for invariants while avoiding redundancy. In particular, by restricting to suitable indecomposable modules associated to each embedding, we obtain bases consisting entirely of interval modules. This approach leads to explicit and computable descriptions of invariants, as well as new bases generalizing the rectangle modules.
@@ 5.1. Interval modules and the intermediate extension functor
It is well known that the simple modules form a basis for the dimension vector, and that the rectangle modules form a basis for the rank invariant. While the latter was established in [BOO24] by an explicit computation, the underlying conceptual reason for the appearance of rectangle modules has remained less clear.
Our aim is to provide a unified explanation of these phenomena using the intermediate extension functor introduced in [AET25]. This functor allows us to associate to each embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$ and each indecomposable $\mathcal{X}$-module $U$ a canonical $\mathcal{P}$-module $\Theta_{f}U$, which behaves well with respect to interval structures. In particular, we show that both simple modules and rectangle modules arise naturally from this construction, corresponding to embeddings of the posets ${\mathcal{X}}_{1}$ and ${\mathcal{X}}_{2}$, respectively.
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$ be the poset embedding given by $fa=(2,1)$, $fb=(1,2)$, $fc=(3,2)$, and $fd=(2,3)$. Consider the sincere interval module $\mathbb{I}_{[\mathcal{X}]}\in\operatorname{\mathrm{mod}}\mathcal{X}$. Since the induction functor $f_{!}$ preserves projective presentations and the coinduction functor $f_{*}$ preserves injective copresentations, we then have
[…]
@@
Let $S\subseteq\mathcal{P}$. The convex hull of $S$ in $\mathcal{P}$ is the set
@@
Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$ be a poset embedding and let $U\in\operatorname{\mathrm{mod}}\mathcal{X}$. By adjunction and the fact that $f^{*}f_{*}=\mathrm{Id}$, there is a natural isomorphism $\operatorname{\mathrm{Hom}}_{\mathcal{P}}(f_{!}U,f_{*}U)\cong\operatorname{\mathrm{End}}_{\mathcal{X}}(U)$. Let $\theta_{f}U\mathrel{\mathop{\ordinarycolon}}f_{!}U\rightarrow f_{*}U$ denote the image of the identity map $1_{U}\in\operatorname{\mathrm{End}}_{\mathcal{X}}(U)$ under this isomorphism. We denote $\Theta_{f}U=\mathrm{Im}(\theta_{f}U)$.
@@
[AET25, Proposition 4.16] Let $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$ be a poset embedding and $\mathbb{I}_{S}\in\operatorname{\mathrm{mod}}\mathcal{X}$ an interval module. Then $\Theta_{f}(\mathbb{I}_{S})\cong\mathbb{I}_{\overline{f(S)}}$.
To make this perspective more concrete, we now examine how the intermediate extension functor behaves in low-dimensional cases. In particular, we compare it with the induction and coinduction functors associated to embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$.
@@
    We consider embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}_{2}\rightarrow\mathcal{P}$, where $\mathcal{X}_{2}=\{1\rightarrow 2\}$. Recall from Example 3.7 that the three indecomposable $\mathcal{X}_{2}$-modules are $\mathbb{I}_{\{1,2\}}=P_{1}=I_{2}$, $\mathbb{I}_{\{1\}}=H_{1,2}=I_{a}$, and $\mathbb{I}_{\{2\}}=P_{2}=C_{1,2}$. For $f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}$, one computes
@@
    In particular, we obtain that $\{\Theta_{f}U\mid f\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}},U\in\operatorname{\mathrm{Ind}}\mathcal{X}\}$ is the set of rectangle modules. Note that there is some redundancy in this collection: for $x\in\mathcal{P}$ we obtain the (simple) rectangle module $\mathbb{I}_{\{x\}}$ once for every embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}_{2}\rightarrow\mathcal{P}$ with either $x=f(1)$ or $x=f(2)$.
@@
    but the invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ does not factor through $\mathbb{Z}^{\mathrm{E}\times\mathrm{I}/\sim_{\Theta}}$. For example, if $\mathcal{P}=\{a\to b\to c\}$, then we have $\Theta_{f}\mathbb{I}_{\{1\}}=\mathbb{I}_{\{a\}}=\Theta_{g}\mathbb{I}_{\{1\}}$ for $\operatorname{\mathrm{Im}}f=\{a,b\}$ and $\operatorname{\mathrm{Im}}g=\{a,c\}$. However, we have for any $\mathcal{P}$-module $M$ with $\operatorname{\mathrm{rk}}M_{a\to b}\neq\operatorname{\mathrm{rk}}M_{a\to c}$ that